# BlueME: Robust Underwater Robot-to-Robot Communication Using Compact Magnetoelectric Antennas

### v2

@@ 6.2 Power Footprint and Robustness
The proposed BlueME system consumes minimal power, with optimal performance achieved at approximately of power consumption (see Fig. 9). Nonlinear effects such as spring-softening in the piezoelectric materials and spring-hardening in the Metglas layers start to become significant at higher drive strengths, leading to shifts and distortions in the resonant frequency with decreased quality factors [66]. In Fig. 9, increased drive power shifts the resonance peak to lower frequencies and distorts the response shape. Shown earlier in Fig. 8, the received signals at various transmitter power levels and distances, a downward frequency shift is also observed. However, since the receiver experiences only the coupled power from the transmitter, these power levels are much lower, resulting in minimal nonlinear distortion at the receiver. Thus, the observed frequency shift primarily reflects the transmitter’s nonlinear behavior combined with the receiver’s nearly constant resonant response. These effects can be exploited to adjust for resonance mismatches between Tx/Rx arrays by varying the drive strength. We also observed that the antenna performance was not significantly affected by obstacles, water turbidity, or multipath propagation, which are major challenges for traditional acoustic and optical modalities.
@@ 6.4 Comparison with Other Underwater Communication Modalities
First, the radio-based RF methods that operate at medium to high-frequency (MF/HF/VHF) exhibit higher data rates, typically up to 1 Mb/s. However, the attenuation in conductive media is much higher, which limits transmission distances to about – m [68, 69, 70, 71]. These higher-frequency systems may also require specialized antenna calibration to address de-tuning effects caused by changing dielectric properties underwater. To this end, acoustic links have been the preferred standard for underwater applications, offering long-range propagation up to km with relatively modest power consumption [72, 73]. Sound waves couple efficiently into the water and travel long distances with lower attenuation than electromagnetic waves [74, 75, 76]. However, acoustic signals experience substantial multipath, Doppler shifts, and boundary reflections, especially in dynamic marine environments. These effects limit typical data throughput to the order of b/s– kb/s and also raise concerns about potential harm to nearby humans and marine fauna if high-intensity acoustic sources are used.
@@
| System | Advantages | Disadvantages |
Examples | |||||||||||
100 kb/s to 1 Mb/s | 1 m to 10 m | 100 to 1k | [68, 69, 70, 71] | |||||||||||
| Acoustic |
10 b/s to 10 kb/s | 100 m to 100 km | 1k to 10k | [72, 73, 74, 75, 76] | |||||||||||
| Optical |
1 Mb/s to 1 Gb/s | 10 m to 100 m | 1k to 10k | [77, 78, 79, 80, 81] | |||||||||||
100 kb/s to 1 Mb/s | 1 m to 100 m | 100 to 1k | [82, 83, 84, 85, 86] | |||||||||||
100 kb/s to 1 Mb/s | 100 cm to 1 m | 10 to 100 | [87, 88, 89, 90, 91, 92] | |||||||||||
• Affected by water conductivity | 1 kb/s to 100 kb/s | 10 m to 1 km | 10 to 100 | This Work |
@@
On the other hand, magnetic-field communication relies on near-field inductive coupling via magnetic coils. These methods are often employed for through-the-earth or short-to-medium-range underwater links and can operate over 1 m to 100 m with data rates up to about 1 Mb/s [82, 83]. Magnetic fields typically penetrate obstacles effectively and offer high reliability where direct line-of-sight may be obstructed. However, large inductors or ferrite materials are often required to generate sufficient field strength, leading to heavier hardware requirements. Accurate alignment between transmitter and receiver coils is also necessary to maximize coupling efficiency [84, 85, 86]. Typical costs for such systems are on the order of $100 to $1 k USD.
@@
Lastly, electric-field communication relies on generating and sensing electric fields in conductive media using capacitive or galvanic coupling [87, 88, 90, 91, 92]. It can be implemented with lightweight and inexpensive electrodes, potentially supporting simultaneous power transfer and data communication. Still, the typical range is limited to about m and there are important safety considerations when using higher voltages around marine life or humans [89].
@@
In comparison, our ME-based radio-frequency (RF) systems in the very-low-frequency (VLF), low-frequency (LF), or medium-frequency (MF) range rely on magnetostrictive-piezoelectric hetero-structures that convert magnetic excitation into electrical signals, and vice versa [11]. One of their primary advantages is immunity to turbidity and obstructions, which often hinder optical and acoustic communication. The longer electromagnetic wavelengths (on the order of tens to hundreds of meters in freshwater) reduce phase interference and Doppler effects, thereby mitigating multipath fading. Although water conductivity does lead to attenuation at increasing distances, the experimentally validated propagation ranges remain practically useful. Fabrication costs for ME-based antennas at the scales used in our experiments can be on the order of a few USD per element, making them relatively cost-effective. Power requirements are low, and the environmental impact is minimal, as these systems generate negligible acoustic noise and avoid light emissions that could disturb marine ecosystems.
@@
Overall, the comparison in Table 1 indicates that no single modality fully addresses all underwater communication requirements. Acoustic systems excel at long-range transmission but are constrained by low data rates. Higher-frequency RF and optical methods can provide larger bandwidths but at severely limited ranges in conductive or turbid environments. By contrast, ME-based RF offers a promising balance of moderate-range coverage, moderate data rates, low power consumption, and low risk to marine life. This trade-off analysis suggests that the BlueME system can occupy an attractive niche in scenarios where turbidity or multipath might compromise other modalities while remaining cost-effective and environmentally benign.
@@ 6.5 Future Potential and Practicalities
In future implementations, it may be desirable to implement frequency shift-keying (FSK) schemes for digital communication. FSK minimizes the impact of amplitude-based nonlinear distortions, allowing for reliable data transmission even in the presence of nonlinearities at higher drive strengths. The ability to pack multiple antennas into arrays provides opportunities for scaling communication capabilities. Bandwidth and power can be increased by adding more antennas, and arrays can be optimized for specific operational frequencies or bandwidth requirements.
@@
7 Conclusions & Future Work
In this paper, we have presented the design, implementation, and experimental validation of a novel ME antenna array system, BlueME, for underwater communication in robotics applications. The BlueME system offers an alternative to traditional acoustic and optical communication methods, addressing challenges in range, efficiency, robustness, and scalability. In testing, the BlueME system achieved reliable communication at distances of with a power consumption of only , indicating its suitability for low-power platforms, such as mobile robots and standalone sensors. Our results also demonstrate that the system remains unaffected by underwater noise, turbidity, and multipath interference.
Future work will focus on optimizing the antenna fabrication process to ensure consistent performance and on tailored modulation schemes that extend the system’s capabilities. This includes enabling applications in multi-robot cooperative tasks such as localization, navigation, and mapping, where reliable, low-latency, and low-power underwater communication is essential.
Acknowledgements

### v4

@@ 5.4 Power Footprint and Robustness
The proposed BlueME system consumes minimal power, with optimal performance achieved at approximately of power consumption (see Fig. 10). Nonlinear effects such as spring-softening in the piezoelectric materials and spring-hardening in the Metglas layers start to become significant at higher drive strengths, leading to shifts and distortions in the resonant frequency with decreased quality factors [67]. In Fig. 10, increased drive power shifts the resonance peak to lower frequencies and distorts the response shape. Shown earlier in Fig. 8, the received signals at various transmitter power levels and distances, a downward frequency shift is also observed. However, since the receiver experiences only the coupled power from the transmitter, these power levels are much lower, resulting in minimal nonlinear distortion at the receiver. Thus, the observed frequency shift primarily reflects the transmitter’s nonlinear behavior combined with the receiver’s nearly constant resonant response. These effects can be exploited to adjust for resonance mismatches between RX/TX arrays by varying the drive strength. We also observed that the antenna performance was not significantly affected by obstacles, water turbidity, or multipath propagation, which are major challenges for traditional acoustic and optical modalities.
@@ 5.5 Channel Capacity
6 Performance Evaluation: Saltwater
Antenna performance was further evaluated in saltwater trials conducted off the west coast of Florida along the Gulf (see Fig. 9). The primary objective was to examine the effect of higher conductivity () on achievable communication range. The experimental setup mirrored that of the freshwater tests, with all equipment housed in watertight enclosures and mounted on the BlueBoat (Tx) and BlueROV2 (Rx). Minor modifications included the integration of a LattePanda single-board computer for autonomous control and the replacement of the original Class-D amplifier with a smaller unit to accommodate spatial constraints. During testing, all systems were powered by onboard batteries and operated fully submerged, except for an Ethernet tether used to stream real-time data from the receiver. For the saltwater trials, the LNA gain was reduced to 40 dB from 60 dB in the freshwater tests to prevent saturation, which lowered the measured noise floor but did not measurably affect the overall SNR.
Autonomous frequency sweeps between to were conducted at one-minute intervals by the transmitter. The BlueBoat carrying the transmitter was allowed to drift freely, with its position tracked using GPS. The receiver was positioned approximately from a stationary pontoon boat and recorded peak-hold FFT measurements each time new GPS coordinates were logged.
An initial transmitter power sweep (shown in Fig. 16) confirmed negligible differences in antenna characteristics relative to the freshwater experiments (shown earlier in Fig. 10), validating the consistency of power characteristics across test conditions. Subsequent experiments involved transmitter sweeps at fixed distances under varying input power levels, as well as tests at fixed transmitter power with varying transmitter–receiver separation. From these measurements, the received SNR was computed, and the corresponding spectral characteristics were analyzed, as presented in Figs. 13 and 14.
6.1 Wavelength and Field Regions
The near-field, transition, and far-field boundaries were estimated using the same method applied in the freshwater analysis. Conductivity at the site was and the water temperature ranged from 20–25∘C [69]. For a center frequency of and seawater conductivity of , the wavelength is calculated as . The corresponding theoretical boundaries for the antenna regions are defined as:
Reactive near-field ():
Radiative near-field ():
Transition region ():
Far-field ():
The measured SNR at distances below approximately exhibited a steeper decay than the nominal – dependence typically associated with reactive or radiative near-field behavior. At larger separations (), the attenuation trend approached an approximate dependence, suggesting that signal propagation may have occurred along paths not confined to the bulk medium, such as along the water–air or water–seabed interfaces [63, 70]. A localized deviation from this overall trend was observed at a transmitter–receiver spacing of (Fig. 13), where the peak received signal level was approximately higher than that measured at . A similar deviation was noted at , where the peak signal level exceeded that at by roughly . A reduction in the measured noise floor was observed during the test, caused by a drop in the LNA supply voltage as its battery approached full discharge. All measurements were performed under identical experimental conditions, with the submerged transmitter naturally drifting and rotating during the tests.
6.2 Radiation Resistance
The same relationship used in the freshwater case can be extended to saltwater by considering the higher conductivity of the medium. Starting from the general dependence
and substituting for , the radiation resistance becomes directly dependent on the conductivity of the surrounding medium. Assuming the frequency, permeability, and relative permittivity remain constant between freshwater and saltwater, this simplifies to
Substituting representative conductivities for freshwater () and saltwater () gives the expected ratio as
This result follows directly from the same scaling principles but reflects the influence of saltwater’s greater conductivity on the radiation resistance of the antenna.
6.3 Power Consumption and Nonlinear Effects
The power consumption characteristics measured during the saltwater tests (Fig. 16) were consistent with those obtained earlier in freshwater (Fig. 10). The same antenna assemblies were used in both experiments, and the response curves exhibited comparable shapes and resonance features. Nonlinear behavior became evident at higher drive levels, with the resonance peak broadening and shifting asymmetrically toward lower frequencies. In the received-signal measurements (Fig. 14), the overall signal amplitude increased with drive strength up to approximately , beyond which additional input power did not produce proportionally higher received levels. A deviation from this trend was observed at an input of , where the peak received signal level exceeded that measured at and .
6.4 Modulation and Practical Considerations
When an ME antenna is driven at high amplitudes, its effective stiffness becomes amplitude dependent, causing the resonance to deviate from the Lorentzian form and exhibit Duffing-type nonlinear behavior [71, 72, 73, 74]. When a modulated signal is applied, these nonlinearities generate harmonic and intermodulation distortion [75, 76]. Amplitude-varying modulation formats are particularly susceptible to spectral regrowth and amplitude-to-phase conversion. Constant-envelope, noncoherent formats such as MSK, CPFSK, and BFSK are therefore preferred for communication; alternative approaches such as direct antenna modulation (DAM) and nonlinear antenna modulation (NAM) have also been explored to intentionally utilize the nonlinear response within the modulation process [77, 78].
ME antennas inherently support multiple resonant modes beyond the fundamental, corresponding to higher-order mechanical and electromechanical harmonics of the piezoelectric element [79]. In the present system, additional resonances appear above approximately 70 kHz and extend to several hundred kHz. Operation outside the fundamental band can excite these higher modes if the transmit spectrum includes harmonic components or spectral leakage from the modulation process. Maintaining spectral confinement helps suppress spurious excitation of these bands, although such measures are unnecessary when operation is restricted to a single frequency range.
The radiated field of an ME antenna originates from the time-varying magnetic flux density produced by magnetostrictive motion, which undergoes a phase transition of nearly through resonance [80]. This frequency-dependent phase transition produces a corresponding frequency-dependent group delay in the radiated signal. The overall group delay of the communication link reflects the combined phase responses of the Tx and Rx antennas. For narrowband operation (), this delay remains effectively constant and waveform distortion is negligible, whereas for wider bandwidths (), phase nonuniformity can distort the transmitted signal unless characterized and compensated at both the transmitter and receiver. Given these antenna characteristics, the desirable properties of a suitable modulation scheme are: (i) constant envelope, to minimize nonlinear conversion; (ii) spectral compactness, to avoid excitation of adjacent resonant modes; and (iii) resilience to phase dispersion, or compatibility with digital equalization schemes.
Binary frequency shift keying (BFSK) is a representative modulation scheme satisfying these requirements when operation is confined to the fundamental resonance band. In BFSK, binary symbols are conveyed by switching the carrier between two frequencies, and , ideally within the resonance bandwidth. Because the envelope remains constant and the frequency deviation can be selected to stay within the array’s linear operating range, BFSK avoids AM-PM conversion and minimizes distortion. Under small-signal conditions where each antenna element operates well below its nonlinearity threshold, theoretical performance can be analyzed using the standard BER–SNR relationship.
Noncoherent BFSK compares the received energy in two frequency bands, so the decision statistic depends only on signal magnitude and is insensitive to carrier phase. This phase independence makes it well-suited to ME antennas, where the radiated and detected phase varies across the resonance bandwidth unless explicitly characterized and compensated at the transmitter and receiver. Tone separation can therefore span a large fraction of the resonant band without affecting detection. The corresponding BER in additive white Gaussian noise (AWGN) [81, 82, 83] is computed as:
The maximum theoretical bit rate for BFSK is determined by the minimum frequency separation required to maintain symbol orthogonality. For noncoherent detection, orthogonality is achieved when the two tones differ by an integer number of cycles over one bit interval , such that . More generally, the energy-per-bit to noise-density ratio can be expressed in terms of the SNR, the bit rate , and the tone spacing as
where is taken as the minimum SNR of the two selected tones. Substituting this into the noncoherent BFSK bit-error rate equation can be reduced as follows:
Given the measured SNR as a function of frequency for various transmission distances (Fig. 13 and Fig. 8), the maximum achievable bit rate for a desired can be estimated by selecting two tones near the resonance frequency that satisfy the BFSK tone-spacing condition and exhibit approximately equal SNR. For a chosen BER threshold, the maximum tone separation that satisfies this criterion can be used to determine , which can then be mapped to a range via the corresponding SNR–distance relationship. This analysis, assuming an uncoded channel, is illustrated in Fig. 15.
7 Analysis & Discussion
@@ 7.1 Comparison with Other Underwater Communication Modalities
| System | Advantages | Disadvantages |
Examples | |||||||||||||
| Acoustic |
10 b/s to 10 kb/s | 100 m to 100 km | 1k to 10k | [84, 85, 86, 87] | |||||||||||||
| Optical |
1 Mb/s to 1 Gb/s | 10 m to 100 m | 1k to 10k | [88, 89, 90, 91, 92] | |||||||||||||
100 kb/s to 1 Mb/s | 1 m to 100 m | 100 to 1k | [93, 94, 95, 96, 97] | |||||||||||||
100 kb/s to 1 Mb/s | 100 cm to 1 m | 10 to 100 | [98, 99, 100, 101, 102, 103] | |||||||||||||
100 kb/s to 1 Mb/s | 1 m to 10 m | 100 to 1k | [104, 105, 106, 107] | |||||||||||||
10 m to 100 m | 100 to 1k | [108, 109] | |||||||||||||
• Affected by water conductivity | 1 kb/s to 100 kb/s | 10 m to 1 km | 10 to 100 |
Acoustic links have been the preferred standard for underwater applications, offering long-range propagation up to km with relatively modest power consumption [84, 85]. Sound waves couple efficiently into the water and travel long distances with lower attenuation than electromagnetic waves [86, 87]. However, acoustic signals experience substantial multipath, Doppler shifts, and boundary reflections, especially in dynamic marine environments. These effects limit typical data throughput to the order of b/s– kb/s and also raise concerns about potential harm to nearby humans and marine fauna if high-intensity acoustic sources are used.
@@
Magnetic-field communication relies on near-field inductive coupling via magnetic coils. These methods are often employed for through-the-earth or short-to-medium-range underwater links and can operate over 1 m to 100 m with data rates up to about 1 Mb/s [93, 94]. Magnetic fields typically penetrate obstacles effectively and offer high reliability where direct line-of-sight may be obstructed. However, large inductors or ferrite materials are often required to generate sufficient field strength, leading to heavier hardware requirements. Accurate alignment between transmitter and receiver coils is also necessary to maximize coupling efficiency [95, 96, 97]. Typical costs for such systems are on the order of $100 to $1 k USD.
@@
Electric-field communication relies on generating and sensing electric fields in conductive media using capacitive or galvanic coupling [98, 99, 101, 102, 103]. It can be implemented with lightweight and inexpensive electrodes, potentially supporting simultaneous power transfer and data communication. Still, the typical range is limited to about m and there are important safety considerations when using higher voltages around marine life or humans [100].
Standard RF methods that operate at medium to high-frequency (MF/HF/VHF) exhibit higher data rates, typically on the order of 1 Mb/s. However, the attenuation in conductive media is much higher, which limits transmission distances to about - m [104, 105, 106, 107]. These higher-frequency systems may also require specialized antenna calibration to address de-tuning effects caused by changing dielectric properties underwater. Recent demonstrations of surface-electromagnetic-wave (SEW)-based RF communication have reported ranges exceeding 100 m and bandwidths of several megahertz by utilizing near-surface propagation along the water-air interface [108, 109]. In these implementations, electromagnetic energy is guided along the boundary layer rather than through the bulk medium, resulting in lower effective attenuation compared to volumetric propagation. This mode of operation represents a distinct configuration from conventional submerged RF systems and is therefore listed as a separate class in Table 1.
@@
In comparison, our ME-based RF systems in the VLF/LF, or medium-frequency (MF) range rely on magnetostrictive-piezoelectric hetero-structures that convert magnetic excitation into electrical signals, and vice versa [11]. One of their primary advantages is immunity to turbidity and obstructions, which often hinder optical and acoustic communication. The longer electromagnetic wavelengths (on the order of tens to hundreds of meters in freshwater) reduce phase interference and Doppler effects, thereby mitigating multipath fading. Although water conductivity does lead to attenuation at increasing distances, the experimentally validated propagation ranges remain practically useful. Fabrication costs for ME-based antennas at the scales used in our experiments can be on the order of a few USD per element, making them relatively cost-effective. Power requirements are low, and the environmental impact is minimal, as these systems generate negligible acoustic noise and avoid light emissions that could disturb marine ecosystems.
@@
Overall, the comparison in Table 1 indicates that no single modality fully addresses all underwater communication requirements. Acoustic systems excel at long-range transmission but are constrained by low data rates. While higher-frequency RF and optical methods can provide larger bandwidths, they operate at severely limited ranges in conductive or turbid waterbody conditions. In contrast, ME-based RF offers a promising balance of moderate-range coverage, moderate data rates, low power consumption, and low risk to marine life. This trade-off analysis suggests that the BlueME system can occupy an attractive niche in scenarios where turbidity or multipath might compromise other modalities while remaining cost-effective and environmentally benign.
@@ 7.2 Future Potential and Practicalities
In future implementations, it may be desirable to implement custom FSK schemes for digital communication. FSK will potentially minimize the impact of amplitude-based nonlinear distortions, allowing for reliable data transmission even in the presence of nonlinearities at higher drive strengths. The ability to pack multiple antennas into arrays provides opportunities for scaling communication capabilities. Bandwidth and power can be increased by adding more antennas, and arrays can be optimized for specific operational frequencies or bandwidth requirements.
@@
8 CONCLUSION
In this paper, we have presented the design, implementation, and experimental validation of a novel ME antenna array system, BlueME, for underwater communication in robotics applications. The BlueME system offers an alternative to traditional acoustic and optical communication methods, addressing challenges in range, efficiency, robustness, and scalability. In testing, the BlueME system achieved reliable communication at distances of over with a power consumption of only in fresh water and over in sea water with a consumption of , and indicating its suitability for low-power platforms, such as mobile robots and standalone sensors. Our results also demonstrate that the system remains unaffected by underwater noise, turbidity, and multipath interference.
Future work will focus on optimizing the antenna fabrication process to ensure consistent performance and on tailored modulation schemes that extend the system’s capabilities. This includes enabling applications in multi-robot cooperative tasks such as localization, navigation, and mapping, where reliable, low-power underwater communication is essential.
ACKNOWLEDGMENT